help

Consider five numerical values, 6.11, 7.04, -6.56,

Consider five numerical values, 6.11, 7.04, -6.56,
Answer #1

mean=(X1+X2+X3+X4+X5)/5=(6.11+7.04+(-6.56)+3.23+14.4)/5=4.844=4.85

OPTION 7 IS CORRECT
Answer #2

mean=(6.11+7.04-6.56+3.23+14.4)/5

=4.85
Answer #3

hi.. answer is 7) 4.85
while calculating mean,,you have to sum up all the numbers and then divide it with 5 and you will get your answer. hope dis was useful :)